OpenStudy (anonymous):

given the equation y-3=1/2(x+6) in point-slope form, identify the equation of the same line in slope-intercept form

OpenStudy (campbell_st):

distribute the right hand side and the add 3 to both sides...

OpenStudy (anonymous):

please solve

OpenStudy (campbell_st):

y - 3 = x/2 + 3 you can do the rest

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

y=1/2x